# Towards Optimal Environmental Policies: Policy Learning under Arbitrary Bipartite Network Interference

### v2

@@ 4.1.1 Outcome Level Data: Medicare Beneficiary Data
The outcome data are derived from Medicare inpatient claims data for all 27,312,190 fee-for-service beneficiaries residing in the U.S. in 2005. The health outcome of interest is the ZIP code-level ischemic heart disease (IHD) inpatient hospitalization rate per 10,000 person-years in 2005 among these Medicare beneficiaries. For each ZIP code, U.S. Census socioeconomic and demographic features (from the 2000 decennial census), meteorological (Kalnay et al., 1996), and smoking rate (Dwyer-Lindgren et al., 2014) covariate data were obtained. Further, due to the relatively small number of power plants in the Western U.S., we remove ZIP codes in the U.S. states of Wyoming, Colorado, Idaho, Utah, Arizona, New Mexico, Nevada, California, Oregon, Montana, and Washington. After initial cleaning, we have $n=26,674$ ZIP codes. The ZIP code covariates used in our analysis, and corresponding descriptive statistics, are given in Table B.2.
The ZIP code-level IHD hospitalization rate per 10,000 person-years is used as the outcome in our statistical models, in order to account for different amounts of person-time of exposure across ZIP codes. Our primary results shown below are based on minimizing rates of hospitalization. The results stated in terms of total counts of hospitalizations are detailed in Section E.3.
@@ 4.1.2 Intervention Level Data: Power Plant Data
Data on scrubber status in the year 2005 and plant characteristics for $J=409$ coal–fired power plants concentrated in the Eastern US, which serve as our intervention units of interest, were obtained from the U.S. Environmental Protection Agency Air Markets Program Database. The plant covariates used in our analyses, along with descriptive statistics, are given in Table B.1. Scrubber installation cost information for power plants that had a scrubber installed by 2021 was obtained from the Energy Information Administration (EIA) website.
4.1.3 Interference Map: HyADS Climate Model
To characterize the interference structure or the bipartite adjacency matrix, we utilize the HyADS model. HyADS is a pollution transport model that yields a unit-less metric quantifying the amount of emissions from an individual power plant that were transported (e.g., by wind) to a particular ZIP code (Henneman et al., 2019). These values, which were calculated for all power plant pairs and ZIP codes in the data, form the elements of $H$.
@@ 4.3 Cost Model Development
In order to find the optimal allocation of scrubbers under cost-constraints, we need to quantify the cost of installing a scrubber at each power plant. Because only a subset of plants has scrubbers installed (and therefore has observed installation cost information), we build a model to impute scrubber installation costs for power plants without scrubbers based on covariates available from our intervention dataset. Since the cost data from the EIA include information from various stacks within a given power plant facility, we define the cost ($c_{j}$ in Equation 4) to be the average cost of installing a scrubber at a given power plant.
The EIA data report scrubber installation costs for 135 of the power plants represented in our data. These 135 samples serves as our training set. We follow a standard machine learning approach to build our cost model. First, we split our training data ($J=135$) into a (sub) training and validation set using an 80/20 split, resulting in ($J=108$) and ($J=27$) respectively. Then, we test three models—namely, random forest, support vector regression, and linear regression—on the validation set. The model with the best performance is selected. Finally, the selected model is then trained on our full training data, and used to generate scrubber installation cost predictions on power plant data with missing cost information ($J=225$). Performance metrics on the validation set, variables’ importance metrics from our final random forest model, and cost predictions are shown in Section D.2. Under our cost model, the total cost for installing a scrubber at all of the 388 power plants is approximately \$43,969,292,000.
@@ 4.4 Policy Analysis
We apply the approaches developed above to (1) quantify the effects of FGD scrubber installation at each power plant on IHD hospitalizations in 2005 and (2) identify optimal scrubber installation policies under cost constraints. For our primary analysis, we use A-Learning to protect against model misspecification.
@@ 4.4.1 TotalEffect Rates Analysis
For interpretability, we build an outcome model and a propensity score model akin to Equation 15 and Equation 16, respectively, but linear in covariates instead of quadratic. Sensitivity analyses with more complex models are reported in Section E.2 and support the conclusions of the linear model with larger IHD reductions found. Upon fitting the A-Learning model, we find that $100\%$ of power plants have estimated protective effects if treated, indicating that scrubber installation would reduce IHD hospitalizations. Graphically Fig. 1 shows a plot of $\widehat{\texttt{TotalEffect}}_{j}$ on the U.S. map and in a histogram. Installing scrubbers at most power plants tend to have small effects, save for a handful of influential power plants in the Midwest. For example, consider the power plant that has the greatest estimated reduction in IHD hospitalization rates if a scrubber in installed, found in Tennessee. If we install a scrubber at this power plant, we would find a 0.61 (408) reduction in IHD hospitalization rate per 10,000 person-years (counts) in 2005 among Medicare beneficiaries. An exploratory analysis testing $H_{0}:\texttt{TotalEffect}_{j}\geq 0,H_{a}:\texttt{TotalEffect}_{j}<0$ based on Theorem 1 yields evidence of statistically significant protective population effects when installing scrubbers at the majority of the power plants for significance level $\alpha=0.05$; Fig. E.1 shows $\sim 95\%$ of p-values are statistically significant. See Supplementary Materials for more information, including p-values and 95% CIs.
Refer to caption Refer to caption
(a) (b)
Figure 1: (a) displays the plot of $\widehat{\texttt{TotalEffect}}_{j}$ on a U.S. map. (b) displays the histogram of $\widehat{\texttt{TotalEffect}}_{j}$.
@@ 4.4.2 Optimal scrubber policies under cost constraints
Because, to our knowledge, policymaking entities have not specified exact budgets for scrubber installations, here we consider various total cost budgets defined as percents (10%, 20%,…,90%) of the estimated cost of installing scrubbers at all power plants used in our analysis (\$43,969,292,000, as reported above). Under each cost budget, the optimal treatment policy, i.e., the set of plants that are selected for scrubber installation, and the subsequent average reduction in IHD hospitalizations per 10,000 person years are shown in Fig. 2. Equivalent analysis for counts are shown in Section E.3. Under a budget constituting 10% of the cost of universal scrubber installations at all US power plants, annual IHD hospitalizations can be reduced by 13,353 (20.66 per 10,000 person-years) via the optimal scrubber installation policy. The plants selected for scrubbers under this budget are largely concentrated in the Ohio River Valley and Midwest. Under a budget constituting 90% of universal installation costs, 31,447 (44.25 per 10,000 person-years) IHD hospitalizations can be avoided annually. As shown in Fig. 2, the health benefits of additional scrubber installations seems to level off for budgets above approximately 60% of universal installation cost, and the biggest reductions occur as the budget increases from 10% – 30% of universal cost.
@@
To investigate this, we map the $\widehat{\texttt{TotalEffect}}_{j}$ alongside the $\widehat{\texttt{BenefitCost}}_{j}$ in Fig. 3. The scale of $\widehat{\texttt{BenefitCost}}_{j}$ is much smaller, so we negate and log transform it for visualization. This enables us to compare how ‘dark’ different power plants are in Fig. 3, to give more insights into the factors influencing selection for treatment in the optimal policies. As we can see, the power plants around the Midwest that have the greatest reductions in $\widehat{\texttt{TotalEffect}}_{j}$ generally have competitive cost-benefit ratios. Interestingly, we can observe that several other power plants in the surrounding regions are also prioritized in the cost-constrained intervention strategies because, although their $\widehat{\texttt{TotalEffect}}_{j}$ is relatively small, they also have cheaper scrubber installation costs, leading to a more competitive $\widehat{\texttt{BenefitCost}}_{j}$.
@@
We end with a comparison of the cost-constrained policies from Fig. 2 to a more intuitive policy that policymakers may carry out under realistic cost constraints. Recall from Section 2.4 that the optimal policy to Equation 4 is obtained from the fractional knapsack solution. Informally, this means that one must: (a) estimate $\texttt{BenefitCost}_{j}$, and (b) treat power plants $j$ with the lowest $\widehat{\texttt{BenefitCost}}_{j}$ (as lower hospitalizations are better) so long as the cost constraint is satisfied. Intuitively, a policymaker may do the same as above, but instead treat power plants with the lowest $\widehat{\texttt{TotalEffect}}_{j}$ (i.e. the greatest reduction in IHD hospitalizations) rather than the lowest $\widehat{\texttt{BenefitCost}}_{j}$; Fig. 4 shows the effect of enacting such a policy. Comparing Fig. 2 to Fig. 4, we find uniformly lower reductions (i.e. greater health benefits) from solving Equation 4 compared to a more intuitive solution.
Refer to caption Refer to caption Refer to caption
(a) 10% Total Cost (b) 20% Total Cost (c) 30% Total Cost
Refer to caption Refer to caption Refer to caption
(a) 40% Total Cost (b) 50% Total Cost (c) 60% Total Cost
Refer to caption Refer to caption Refer to caption
(a) 70% Total Cost (b) 80% Total Cost (c) 90% Total Cost
Figure 2: Grid with the reduction of IHD Hospitalizations/10,000 Person Years, varying the spending from 10%-90% of budget
Refer to caption Refer to caption
(a) (b)
Figure 3: (a) displays the plot of $-\widehat{\texttt{TotalEffect}}_{j}$ on a U.S. map, colored by intensity of $\widehat{\texttt{TotalEffect}}_{j}$. (b) displays an analagous plot for $-\log(-\widehat{\texttt{BenefitCost}}_{j})$.
Refer to caption Refer to caption Refer to caption
(a) 10% Total Cost (b) 20% Total Cost (c) 30% Total Cost
Refer to caption Refer to caption Refer to caption
(a) 40% Total Cost (b) 50% Total Cost (c) 60% Total Cost
Refer to caption Refer to caption Refer to caption
(a) 70% Total Cost (b) 80% Total Cost (c) 90% Total Cost
Figure 4: Grid with the reduction of IHD Hospitalizations/10,000 Person Years, varying the spending from 10%-90% of budget, optimizing based on $\widehat{\texttt{TotalEffect}}$ rather than $\widehat{\texttt{BenefitCost}}$.
@@ 5 Discussion
Understanding how to optimally assign an intervention is critically important to optimize policy effects while keeping the costs fixed. In this paper, we proposed a new methodology to produce optimal policies in the context of bipartite network interference. Using our proposed methodology, we carry out the first statistical policy learning analyses for air pollution policy, finding up to 31,906 (44.51 per 10,000 person-years) IHD hospitalizations can be avoided annually among Medicare fee-for-service beneficiaries. This analysis supports previous findings that reducing pollutants from large-scale emissions reduces public health burden (Henneman et al., 2023). The analysis of $\widehat{\texttt{BenefitCost}}$ sheds light onto how nuanced this analysis can be under realistic cost constraints. Different budgets and policy choices can yield variable effects on public health impact.
@@
From a methodological standpoint, we proposed the first BNI policy learning methodology based on Q-Learning and A-Learning under arbitrary BNI and validated our proposed methods in an extensive simulation study that closely mimics our real world setting. Our approach does not require clustering assumptions or defining direct and spillover estimands as commonly done in interference literature; such assumptions are untenable in this problem due to air pollution transport. The proposed methodologies hold general applicability for future large environmental policy studies, as it can encompass variants of clustered BNI.
@@
There are several limitations to note. We conduct our primary analysis assuming a linear exposure mapping and that HyADS is a primary effect modifier for power plant interventions. Such an assumption enables computationally feasible modeling and aligns with domain knowledge, but it is untestable in practice. Methodological limitations of this work are assuming costs $c_{1}\dots c_{J}$ are known and the interference map $\mathbf{H}$ is known (as in practice HyADS is an estimation method). Regardless, assuming that $\mathbf{H}$ is known, is a common assumption made in bipartite literature (see, e.g., Papadogeorgou et al., 2019; Doudchenko et al., 2020; Chen, Bargagli-Stoffi, Kim & Nethery, 2024; Chen, Bargagli-Stoffi, Kim, Henneman & Nethery, 2024). Future work should address the limitations above. Further, given recent evidence of disproportionate burden to marginalized subgroups in air pollution regulatory policy (Zeger et al., 2008; Hajat et al., 2015; Kioumourtzoglou et al., 2015; Di et al., 2017; Jbaily et al., 2022; Josey et al., 2023; Chen, Bargagli-Stoffi, Kim & Nethery, 2024), extending this work to ensure ‘fair’ policy learning (Viviano & Bradic, 2024) stands as an important future direction.
Funding.
@@
This work was funded by the Health Effects Institute award 4998-RFA22-2/24 and NIH grants T32ES714242, K01ES032458.
@@
1Department of Biostatistics, Harvard T.H. Chan School of Public Health (Boston, MA)
2Department of Biostatistics, UCLA Fielding School of Public Health (Los Angeles, CA)
Appendix A Proofs
The results below employ the same strategy used in Su et al. (2019). However, we require results in a setting where there is arbitrary bipartite network interference. Since the setup is unique, we explicitly write out the results.
A.1 Proof of Proposition 2.1
$\mathbb{E}[Y_{i}(\bar{a}_{i})]$ $=\mathbb{E}[\mathbb{E}[Y_{i}(\bar{a}_{i})\mid\mathbf{X}^{out}_{i},\mathbf{\mathbb{X}}_{1:J}]]$
=(A3)${{=}}\mathbb{E}[\mathbb{E}[Y_{i}(\bar{a}_{i})\mid\mathbf{X}^{out}_{i},\mathbf{\mathbb{X}}_{1:J},\bar{A}_{i}]]$
=(A4)${{=}}\mathbb{E}[\mathbb{E}[Y_{i}(\bar{a}_{i})\mid\mathbf{X}^{out}_{i},\bar{A}_{i}]]$
=(A1)${{=}}\mathbb{E}[\mathbb{E}[Y_{i}\mid\mathbf{X}^{out}_{i},\bar{A}_{i}]]$
A.2 Proof of Q-Learning Inference
Similar to the standard assumptions for M-estimation (Stefanski & Boos, 2002), the following are required for statistical inference of $\mathbb{E}[Y_{i}(\bar{a}_{i};\boldsymbol{\theta}_{0})]$ under Q-Learning:
@@ A.2 Proof of Q-Learning Inference
  1. (A7)
    The outcome covariates, intervention covariates, and interference map are bounded. $||\mathbf{X}^{out}_{i}||_{\infty}<\infty$, $||\mathbf{X}^{int}_{j}||_{\infty}<\infty,||\mathbf{H}||<\infty$
  2. (A8)
    Parameters $\boldsymbol{\theta}$ lie in a compact space $\mathbf{\boldsymbol{\theta}}$ with $\boldsymbol{\theta}_{0}$ as an inner point.
  3. (A9)
@@
Using Taylor’s theorem as in Section 2 of (Stefanski & Boos, 2002):
@@
[…]
Consistent variance estimation follows by $\hat{\Sigma}=\Sigma(\hat{\boldsymbol{\theta}})=\frac{1}{n}\Sigma_{d}^{-1}(\hat{\boldsymbol{\theta}})\Sigma_{\phi}(\hat{\boldsymbol{\theta}})(\Sigma_{d}^{-1}(\hat{\boldsymbol{\theta}}))^{\top}$ for $\Sigma_{d}(\boldsymbol{\theta})=\frac{1}{n}d_{i}(\boldsymbol{\theta})d_{i}(\boldsymbol{\theta})^{\top}$. ∎
A.3 Proof of Lemma 1
@@
$\mathbb{E}[\Phi_{n}(\boldsymbol{\beta}_{0},\boldsymbol{\alpha}_{0},\boldsymbol{\gamma}_{0})]$ =(A1)${{=}}\mathbb{E}[\mathbb{E}[\Phi_{n}(\boldsymbol{\beta}_{0},\boldsymbol{\alpha}_{0},\boldsymbol{\gamma}_{0})\mid\mathbf{X}^{int}_{j},\mathbf{X}^{out}_{i},Y_{i}(\bar{a}_{i})]]$
@@
$\mathbb{E}[\mathbb{E}[\phi_{i}(\boldsymbol{\beta}_{0},\boldsymbol{\alpha}_{0},\boldsymbol{\gamma}_{0})$ $\mid\mathbf{X}^{int}_{j},\mathbf{X}^{out}_{i},Y_{i}(\bar{a}_{i})]]$
@@
$\mathbb{E}[\Phi_{n}(\boldsymbol{\beta}_{0},\boldsymbol{\alpha}_{0},\boldsymbol{\gamma}_{0})]$ $=\mathbb{E}[\mathbb{E}[\Phi_{n}(\boldsymbol{\beta}_{0},\boldsymbol{\alpha}_{0},\boldsymbol{\gamma}_{0})\mid\mathbf{X}^{out}_{i},\mathbf{\mathbb{X}}_{1:J}]]$
=(A3)${{=}}\mathbb{E}[\mathbb{E}[\Phi_{n}(\boldsymbol{\beta}_{0},\boldsymbol{\alpha}_{0},\boldsymbol{\gamma}_{0})\mid\mathbf{X}^{out}_{i},\mathbf{\mathbb{X}}_{1:J},\bar{A}_{i}]]$
=(A4)${{=}}\mathbb{E}[\mathbb{E}[\Phi_{n}(\boldsymbol{\beta}_{0},\boldsymbol{\alpha}_{0},\boldsymbol{\gamma}_{0})\mid\mathbf{X}^{out}_{i},\bar{A}_{i}]]$
@@
$\mathbb{E}[\mathbb{E}[\phi_{ni}(\boldsymbol{\beta}_{0},\boldsymbol{\alpha}_{0},\boldsymbol{\gamma}_{0})$ $\mid\mathbf{X}^{out}_{i},\bar{A}_{i}]]$

### v3

@@ 4.1.1 Outcome Level Data: Medicare Beneficiary Data
The outcome data are derived from Medicare inpatient claims data for all 27,312,190 fee-for-service beneficiaries residing in the U.S. in 2005. The health outcome of interest is the ZIP code-level ischemic heart disease (IHD) inpatient hospitalization rate per 10,000 person-years in 2005 among these Medicare beneficiaries. For each ZIP code, U.S. Census socioeconomic and demographic features (from the 2000 decennial census), meteorological (Kalnay et al., 1996), and smoking rate (Dwyer-Lindgren et al., 2014) covariate data were obtained. Further, due to the relatively small number of power plants in the Western U.S., we remove ZIP codes in the U.S. states of Wyoming, Colorado, Idaho, Utah, Arizona, New Mexico, Nevada, California, Oregon, Montana, and Washington. After initial cleaning, we have $n=26,674$ ZIP codes. The ZIP code covariates used in our analysis, and corresponding descriptive statistics, are given in Table Appendix K.2.
The ZIP code-level IHD hospitalization rate per 10,000 person-years is used as the outcome in our statistical models, in order to account for different amounts of person-time of exposure across ZIP codes. Our primary results shown below are based on minimizing rates of hospitalization. The results stated in terms of total counts of hospitalizations are detailed in Section Appendix N.3.
@@ 4.1.2 Intervention Level Data: Power Plant Data
Data on scrubber status in the year 2005 and plant characteristics for $J=409$ coal–fired power plants concentrated in the Eastern US, which serve as our intervention units of interest, were obtained from the U.S. Environmental Protection Agency Air Markets Program Database. The plant covariates used in our analyses, along with descriptive statistics, are given in Table Appendix K.1. Scrubber installation cost information for power plants that had a scrubber installed by 2021 was obtained from the Energy Information Administration (EIA) website.
Interference Map: HyADS Climate Model
To characterize the interference structure or the bipartite adjacency matrix, we utilize the HyADS model. HyADS is a pollution transport model that yields a unit-less metric quantifying the amount of emissions from an individual power plant that were transported (e.g., by wind) to a particular ZIP code (Henneman et al., 2019). HyADS is based on highly accurate meteorological measures from the HYSPLIT model by the National Oceanic and Atmospheric Administration (NOAA) Air Resources Laboratory. These values, which were calculated for all power plant pairs and ZIP codes in the data, form the elements of $H$.
@@ 4.3 Cost Model Development
In order to find the optimal allocation of scrubbers under cost-constraints, we need to quantify the cost of installing a scrubber at each power plant. Because only a subset of plants has scrubbers installed (and therefore has observed installation cost information), we build a model to impute scrubber installation costs for power plants without scrubbers based on covariates available from our intervention dataset. Since the cost data from the EIA include information from various stacks within a given power plant facility, we define the cost ($c_{j}$ in Eq. 4) to be the average cost of installing a scrubber at a given power plant.
The EIA data report scrubber installation costs for 135 of the power plants represented in our data. These 135 samples serves as our training set. We follow a standard machine learning approach to build our cost model. First, we split our training data ($J=135$) into a (sub) training and validation set using an 80/20 split, resulting in ($J=108$) and ($J=27$) respectively. Then, we test three models—namely, random forest, support vector regression, and linear regression—on the validation set. The model with the best performance is selected. Finally, the selected model is then trained on our full training data, and used to generate scrubber installation cost predictions on power plant data with missing cost information ($J=225$). Performance metrics on the validation set, variables’ importance metrics from our final random forest model, and cost predictions are shown in Section Appendix M.2. Under our cost model, the total cost for installing a scrubber at all of the 388 power plants is approximately \$43,969,292,000.
We clarify that the $J=135$ power plants considered in our training set above represents the number of treated power plants by 2021. In contrast, the number of treated power plants in 2005 is 73 (see Table Appendix K.1). When building our cost model, we utilize all power plants treated with scrubbers by 2021 to build the strongest cost model we can.
@@ 4.4 Policy Analysis
[…]
@@ 4.4.1 TotalEffect Rates Analysis
For interpretability, we build an outcome model and a propensity score model akin to that specified in Section 3, respectively, but linear in covariates instead of quadratic. We additionally augment $\mathbf{X}^{int}_{j}$ with weighted summaries of outcome level covariates given by $\mathbf{X}^{out}_{j}=\frac{1}{n}\sum_{i=1}^{n}H_{ij}\mathbf{X}^{out}_{i}$ since outcome covariates and the interference map may impact treatment decisions. Sensitivity analyses with more complex models are reported in Appendix Appendix N.2. Upon fitting the A-Learning model, we find that $96\%$ of power plants have estimated protective effects if treated, indicating that scrubber installation would reduce IHD hospitalizations. Graphically Fig. 1 shows a plot of $\widehat{\texttt{TotalEffect}}_{j}$ on the U.S. map and in a histogram. Installing scrubbers at most power plants tend to have small effects, save for a handful of influential power plants in the Midwest. For example, consider the power plant that has the greatest estimated reduction in IHD hospitalization rates if a scrubber in installed, found in Tennessee. If we install a scrubber at this power plant, we would find a .69 (448) reduction in IHD hospitalization rate per 10,000 person-years (counts) in 2005 among Medicare beneficiaries.
When determining which power plants to implement interventions on, policymakers may wish to incorporate statistical uncertainty into the decision-making process. Thus, to enhance the practical utility and policy-relevance of the proposed methodologies, we perform an exploratory analysis testing $H_{0}:\texttt{TotalEffect}_{j}\geq 0,H_{a}:\texttt{TotalEffect}_{j}<0$ based on Theorem 3. The results are shown in Fig. Appendix N.1. Approximately $70\%$ of p-values are statistically significant, yielding evidence of statistically significant protective population effects when installing scrubbers at the majority of the power plants for significance level $\alpha=0.05$.
@@ 4.4.2 Optimal scrubber policies under cost constraints
Because, to our knowledge, policymaking entities have not specified exact budgets for scrubber installations, here we consider various total cost budgets defined as percents (10%, 20%,…,90%) of the estimated cost of installing scrubbers at all power plants used in our analysis (\$43,969,292,000, as reported above). Under each cost budget, the optimal treatment policy, i.e., the set of plants that are selected for scrubber installation, and the subsequent average reduction in IHD hospitalizations per 10,000 person years are shown in Fig. 2. Equivalent analysis for counts are shown in Appendix Appendix N.3. Under a budget constituting 10% of the cost of universal scrubber installations at all US power plants, annual IHD hospitalizations can be reduced by 15,025 (23.37 per 10,000 person-years) via the optimal scrubber installation policy. The plants selected for scrubbers under this budget are largely concentrated in the Ohio River Valley and Midwest. Under a budget constituting 90% of universal installation costs, 37,928 (55.00 per 10,000 person-years) IHD hospitalizations can be avoided annually. As shown in Fig. 2, the health benefits of additional scrubber installations seems to level off for budgets above approximately 60% of universal installation cost, and the biggest reductions occur as the budget increases from 10% – 30% of universal cost.
@@
To investigate this, we map the $\widehat{\texttt{TotalEffect}}_{j}$ alongside the $\widehat{\texttt{BenefitCost}}_{j}$ in Fig. 3. The scale of $\widehat{\texttt{BenefitCost}}_{j}$ is much smaller, so we negate and log transform it for visualization purposes. This enables us to compare how ‘dark’ different power plants are in Fig. 3, to give more insights into the factors influencing selection for treatment in the optimal policies. As we can see, the power plants around the Midwest that have the greatest reductions in $\widehat{\texttt{TotalEffect}}_{j}$ generally have competitive cost-benefit ratios. Interestingly, we can observe that several other power plants in the surrounding regions are also prioritized in the cost-constrained intervention strategies because, although their $\widehat{\texttt{TotalEffect}}_{j}$ is relatively small, they also have cheaper scrubber installation costs, leading to a more competitive $\widehat{\texttt{BenefitCost}}_{j}$.
@@
In Section Appendix N.5 of the Appendix, we directly compare the factual scrubber policy in 2005 with the optimal policy of the same cost, finding that the optimal policy would have treated a larger number of power plants and more than tripled the reductions in IHD hospitalization rates relative to the factual policy. We end with a comparison of the cost-constrained policies from Fig. 2 to a more intuitive policy that policymakers may carry out under realistic cost constraints. Recall from Section 2.4 that the optimal policy to Eq. 4 is obtained from the fractional knapsack solution. Informally, this means that one must: (a) estimate $\texttt{BenefitCost}_{j}$, and (b) treat power plants $j$ with the lowest $\widehat{\texttt{BenefitCost}}_{j}$ (as lower hospitalizations are better) so long as the cost constraint is satisfied. Intuitively, a policymaker may do the same as above, but instead treat power plants with the lowest $\widehat{\texttt{TotalEffect}}_{j}$ (i.e. the greatest reduction in IHD hospitalizations) rather than the lowest $\widehat{\texttt{BenefitCost}}_{j}$; Fig. 4 shows the effect of enacting such a policy. Comparing Fig. 2 to Fig. 4, we find uniformly lower reductions (i.e. greater health benefits) from solving Eq. 4 compared to a more intuitive solution.
In Section Appendix N.6, we check the sensitivity of our findings to potential violations to the temporal ordering between scrubber installations and measured outcomes in 2005. There are only 4 of 388 (1%) of power plants that may violate this, and results are nearly identical whether we include or exclude these plants.
@@ 5 Discussion
Understanding how to optimally assign an intervention is critically important to optimize policy effects while keeping the costs fixed. In this paper, we proposed a new methodology to produce optimal policies in the context of bipartite network interference. Using our proposed methodology, we carry out the first statistical policy learning analyses for air pollution policy, finding up to 38,372 (55.30 per 10,000 person-years) IHD hospitalizations can be avoided annually among Medicare fee-for-service beneficiaries. This analysis supports previous findings that reducing pollutants from large-scale emissions reduces public health burden (Henneman et al., 2023). The analysis of $\widehat{\texttt{BenefitCost}}$ sheds light onto how nuanced this analysis can be under realistic cost constraints. Different budgets and policy choices can yield variable public health impacts. We consider a clean-slate analysis in which none of the power plants have scrubbers installed in order to provide insights into the potential benefits that could be experienced. In the appendix, we design optimal policies that build upon the factual 2005 scrubber landscape rather than a clean-slate. The contrast between the two approaches are shown in Table Appendix N.2, providing insight into how efficient the factual policy may be. If we were interested in decision-making on future power plants, we can leverage our rule from Q-Learning or A-Learning. The details of this procedure are outlined in Appendix Appendix O.
@@
From a methodological standpoint, we proposed the first BNI policy learning methodology based on Q- and A-Learning under arbitrary BNI and validated our proposed methods in an extensive simulation study that closely mimics our real world setting. Our approach does not require clustering assumptions or defining direct and spillover estimands as commonly done in interference literature; such assumptions are untenable in this problem due to air pollution transport. The proposed methodologies hold general applicability for large environmental policy studies.
@@
There are several limitations to note. We conduct our primary analysis assuming a linear exposure mapping and that HyADS is a primary effect modifier for power plant interventions. Such an assumption enables computationally feasible modeling and aligns with domain knowledge, but it is untestable in practice. Methodological limitations of this work are assuming costs $c_{1}\dots c_{J}$ are known. Future work should address the limitations above. Further, given recent evidence of disproportionate burden to marginalized subgroups in air pollution regulatory policy (Zeger et al., 2008; Hajat et al., 2015; Kioumourtzoglou et al., 2015; Di et al., 2017; Jbaily et al., 2022; Josey et al., 2023; Chen et al., 2024b), extending this work to ensure ‘fair’ policy learning (Viviano and Bradic, 2024) stands as an important future direction.
6 Competing interests
No competing interest is declared.
7 Author contributions statement
R.C.K., F.J.B.S., K.L.C., and R.C.N., contributed to writing and conception of the manuscript. R.C.K. and K.L.C. contributed to performing the analysis.
8 Acknowledgments
@@
This work was funded by the Health Effects Institute award 4998-RFA22-2/24 and NIH grants T32ES714242, K01ES032458. No conflicts of interest to declare.
9 Supplementary Material
Information about where the data are obtained from, simulation code, and analysis code are available at https://github.com/NSAPH-Projects/emissions-ihd-bni_optimal_tx. Supplementary material is available online at http://biostatistics.oxfordjournals.org.
@@ 9 Supplementary Material
1Department of Biostatistics, Harvard T.H. Chan School of Public Health, 677 Huntington Ave, 02215, MA, USA
2Department of Biostatistics, UCLA Fielding School of Public Health, 650 Charles E Young Dr S, 90095, CA, USA
Appendix J Proofs
The inferential results below employ the strategy from Su et al. [2019] and the regret bounds employ the strategy from Kuznetsov and Mohri [2017]. However, we require results in a setting where there is arbitrary bipartite network interference. Since the setup is unique, we explicitly write out the results.
Appendix J.1 Proof of Proposition 1
$\mathbb{E}[Y_{i}(\bar{a}_{i})]$ $=\mathbb{E}[\mathbb{E}[Y_{i}(\bar{a}_{i})\mid\mathbf{X}^{out}_{i},\mathbf{\mathbb{X}}_{1:J},\mathbf{H}_{i}]]$
$\stackrel{{\scriptstyle(\textbf{I3})}}{{=}}\mathbb{E}[\mathbb{E}[Y_{i}(\bar{a}_{i})\mid\mathbf{X}^{out}_{i},\mathbf{\mathbb{X}}_{1:J},\mathbf{H}_{i},\bar{A}_{i}=\bar{a}_{i}]]$
$\stackrel{{\scriptstyle(\textbf{I4})}}{{=}}\mathbb{E}[\mathbb{E}[Y_{i}(\bar{a}_{i})\mid\mathbf{X}^{out}_{i},\mathbf{H}_{i},\bar{A}_{i}=\bar{a}_{i}]]$
$\stackrel{{\scriptstyle(\textbf{I1})}}{{=}}\mathbb{E}[\mathbb{E}[Y_{i}\mid\mathbf{X}^{out}_{i},\mathbf{H}_{i},\bar{A}_{i}=\bar{a}_{i}]]$
Appendix J.2 Proof of Q-Learning Inference
@@
  1. (R1)
[…]
  2. (R2)
    Parameters $\boldsymbol{\theta}$ lie in a compact space $\mathbf{\Theta}$ with $\boldsymbol{\theta}_{0}$ as an inner point.
  3. (R3)
@@
Using Taylor’s theorem:
@@
Then invoking 3 and 4, by the multiplier central limit theorem Theorem 6, we find that $\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\phi(Y_{i};\boldsymbol{\theta}_{0})\xrightarrow{d}N(0,\Sigma_{\phi}(\boldsymbol{\theta}_{0}))$ where $\Sigma_{\phi}=\frac{1}{n}\phi(Y_{i};\boldsymbol{\theta})\phi(Y_{i};\boldsymbol{\theta})^{\top}=\frac{1}{n}\sum_{i=1}^{n}d_{i}(\boldsymbol{\theta})d_{i}(\boldsymbol{\theta})^{\top}(Y_{i}-\mu(\boldsymbol{\theta}))^{2}$.
Consistent variance estimation follows by $)\hat{\Sigma}=\Sigma(\hat{\boldsymbol{\theta}})=\frac{1}{n}\Sigma_{d}^{-1}(\hat{\boldsymbol{\theta}})\Sigma_{\phi}(\hat{\boldsymbol{\theta}})(\Sigma_{d}^{-1}(\hat{\boldsymbol{\theta}}$$)^{\top}$ for $\Sigma_{d}(\boldsymbol{\theta})=\frac{1}{n}d_{i}(\boldsymbol{\theta})d_{i}(\boldsymbol{\theta})^{\top}$. ∎
Appendix J.3 Proof of Lemma 1
@@
[…]
@@
$\mathbb{E}[\mathbb{E}[\phi_{i}(\boldsymbol{\beta}_{0},\boldsymbol{\alpha}_{0},\boldsymbol{\gamma}_{0})$ $\mid\mathbf{X}^{int}_{j},\mathbf{X}^{out}_{i},\mathbf{H}_{i},Y_{i}(\bar{a}_{i})]]$
@@
$\mathbb{E}[\Phi_{n}(\boldsymbol{\beta}_{0},\boldsymbol{\alpha}_{0},\boldsymbol{\gamma}_{0})]$ $=\mathbb{E}[\mathbb{E}[\Phi_{n}(\boldsymbol{\beta}_{0},\boldsymbol{\alpha}_{0},\boldsymbol{\gamma}_{0})\mid\mathbf{X}^{out}_{i},\mathbf{\mathbb{X}}_{1:J},\mathbf{H}_{i}]]$
$\stackrel{{\scriptstyle(\textbf{I3})}}{{=}}\mathbb{E}[\mathbb{E}[\Phi_{n}(\boldsymbol{\beta}_{0},\boldsymbol{\alpha}_{0},\boldsymbol{\gamma}_{0})\mid\mathbf{X}^{out}_{i},\mathbf{\mathbb{X}}_{1:J},\mathbf{H}_{i},\bar{A}_{i}=\bar{a}_{i}]]$
$\stackrel{{\scriptstyle(\textbf{I4})}}{{=}}\mathbb{E}[\mathbb{E}[\Phi_{n}(\boldsymbol{\beta}_{0},\boldsymbol{\alpha}_{0},\boldsymbol{\gamma}_{0})\mid\mathbf{X}^{out}_{i},\mathbf{H}_{i},\bar{A}_{i}=\bar{a}_{i}]]$
@@
$\mathbb{E}[\mathbb{E}[\phi_{ni}(\boldsymbol{\beta}_{0},\boldsymbol{\alpha}_{0},\boldsymbol{\gamma}_{0})$ $\mid\mathbf{X}^{out}_{i},\mathbf{H}_{i},\bar{A}_{i}=\bar{a}_{i}]]$